# MultiOCR-QA: Dataset for Evaluating Robustness of LLMs in Question Answering on Multilingual OCR Texts

### v1

@@ 4.1. Quantifying and Filtering OCR Noise
Outlier Detection: To ensure a reliable analysis, we applied the Interquartile Range (IQR) method to detect and remove outliers in CER values. Outliers were defined as CER values below Q1 - 1.5 * IQR or above Q3 + 1.5 * IQR, where Q1 and Q3 represent the 33th and 66th percentiles of the CER distribution, respectively, and IQR is the difference between Q3 and Q1. This filtering resulted in the removal of 1,056 English, 351 French, and 2,423 German paragraphs.
@@
[…]
@@
Mean Substitutions 74.66 31.51 240.23
Median Substitutions 61.00 21.00 230.50
Median Deletions 11.74 41.09 85.09
Median Deletions 7.00 25.00 39.00
Median Insertions 39.02 20.90 82.60
Median Insertions 30.00 11.00 80.00
4.2. Human Evaluation
To assess the quality of the MultiOCR-QA dataset, we also conducted a manual evaluation study. A total of 100 question-answer pairs were randomly selected for evaluation from each language. We asked native speakers for each language, to rate the questions. The evaluation followed a five-point rating scale (1 to 5), where 1 indicated very poor and 5 indicated very good. The assessment was based on four key criteria: (1) Question Readability, assessing grammatical correctness and fluency; (2) Answer Readability, ensuring coherence and grammatical correctness; (3) Relevance, verifying whether the generated question aligns with the passage; and (4) Answer correctness, confirming the accuracy and completeness of the answer.
Each question was rated by an annotator, and the final scores for each criterion were averaged to obtain an overall evaluation. The results of the human evaluation are presented in Table 3. The evaluation showed high scores for both question and answer readability for the English and French language. The scores for relevance and answer correctness were moderate for these languages, indicating that most of the questions generated were contextually appropriate and that the answers provided were generally correct. But the results for German language were very low depicting that the CorrectedOCR text has a huge number of OCR errors so the errors are carried out further in the generated dataset. Such levels of errors are common when digitizing old historical texts due to the physical damage, irregular fonts, or faded ink of original document (Jatowt et al., 2019).
Table 3. Human evaluation results of MultiOCR-QA dataset.
Language Question Readability Answer Readability Relevance Answer Correctness
English 4.13 4.03 3.33 3.75
French 4.16 4.14 3.75 3.30
German 2.80 1.92 1.64 1.59
@@ 5. Experiments and Results
In this section, we conduct a comprehensive analysis of MultiOCR-QA from several perspectives. First, we evaluate the performance of MultiOCR-QA across various LLMs, comparing different model families and sizes to assess their effectiveness in handling OCR-generated text. Second, we investigate the impact of OCR errors on QA performance, focusing on three primary error types: insertion errors, deletion errors, and substitution errors. By analyzing these aspects, we aim to quantify the robustness of LLMs in processing noisy OCR text and provide insights into the challenges associated with question-answering on historical documents.
@@ 5.2. Experimental Results
Tables 4, 5, and 6 present the impact of OCR errors on the performance of various LLMs in question-answering tasks for English, French, and German texts.
Table 4 presents the performance of LLMs on English text using both CorrectedOCR (CP) and RawOCR (RP) paragraphs. Across all models, the transition from CP to RP negatively impacts performance. The best-performing model on CP is Gemma-2 27B with a BERTScore of 62.90, followed by LLaMA-3.3-70B at 60.71. When switching to RP, Gemma-2 still achieves the highest BERTScore 59.70, but with a 5.09% drop, highlighting its robustness. The lowest-performing model is Mixtral 8×22B, showing the most significant impact of OCR errors. Qwen-2.5 72B model achieves the best performance in CP for the Contains metric of 59.95, indicating its strong retrieval ability in clean text. However, it experiences a 9.48% drop in RP, suggesting moderate sensitivity to OCR errors. The lowest performing model in RP is again Mixtral 8×22B for Contains metric which sees an 11.26% decrease, indicating its greater vulnerability to noisy text. The Exact Match (EM) metric shows the steepest decline across models, emphasizing that OCR errors severely impact the models’ ability to generate precise answers.
Table 5 reports the performance of LLMs on French text, again comparing CorrectedOCR and RawOCR paragraphs. The same as in English, the models perform better with CorrectedOCR, confirming and quantifying the negative influence of OCR errors on the QA accuracy. Gemma-2 27B model achieves the highest BERTScore of 76.51 on CP, demonstrating its strong ability to capture semantic similarity. Despite the 1.46% drop, it still maintains the highest performance with a BERTScore of 75.39 on RP, indicating robustness to OCR noise. Mixtral 8×22B model shows the smallest drop of 1.05%, but its overall score remains lower than the ones for the other models. For Contains, Qwen-2.5 72B achieves the highest score on CP 57.55, highlighting superior retrieval performance on clean text. However, it experiences a 19.90% drop in RP, reinforcing its vulnerability to OCR errors. LLLaMA-3.1 8B also struggles, with a 19.63% decline, showing difficulty in retrieving spans from noisy text. In terms of EM, Gemma-2 27B outperforms all models with an EM of 17.46 on CP. Although it drops by 15.23%, it still maintains the best EM value 14.80 in RP. Mixtral 8×22B model is the most affected in EM, dropping by 27.99%, suggesting that OCR noise drastically reduces its accuracy in generating exact answers.
@@
Table 6 examines the performance of the model in the German text, revealing the most significant performance decline among the three languages. Unlike English and French, German exhibits the largest performance drop due to its lower OCR quality, and the models struggle more with its linguistic structure. Among the models evaluated, Gemma-2 27B achieves the highest BERTScore of 67.07 on CP, confirming its strong ability to capture semantic similarity in clean text. It also maintains the best performance on RP (63.78); however, it still experiences a 4.91% decrease, highlighting the adverse effects of OCR errors. In contrast, LLaMA-3.1 8B and LLaMA-3.3 70B show the biggest BERTScore drop (8.74% and 8.43% respectively), suggesting that these models struggle more with OCR noise. For the Contains metric, LLaMA-3.3 70B achieves the highest Contains score on CP, making it the most effective for retrieving relevant information in clean text. However, all models suffer a severe drop in Contains when moving to RP, with reductions exceeding 65%. Qwen-2.5 7B has the worst drop 66.40%, indicating that it faces challenges to retrieve information from noisy text. In terms of EM, Gemma-2 27B achieves the highest EM score on CP 2.69, while Mixtral 8×22B has the lowest EM 0.078. EM scores drop drastically across all models, the largest decrease is 79.04% for Gemma-2, reinforcing that word-level distortions from OCR errors make exact answer matching nearly impossible. The Mixtral model drops 72.74% in EM, making it highly unreliable for exact answers in noisy OCR text.
Summary of findings: OCR errors consistently degrade the performance of the models in English, French and German texts, resulting in maximum drop of 5.09%, 1.75%, and 8.74% in BERTScore, respectively. The most severe impact was observed in German due to the lower quality of the OCR and the complex linguistic structure. While larger models like Gemma-2 27B and LLaMA-3.3 70B demonstrate greater resilience, all models suffer substantial declines in Contains and Exact Match (EM) metrics, highlighting their weakness in retrieving and generating precise answers from noisy text. Gemma-2 27B consistently outperforms others, maintaining the highest BERTScore and EM across all languages, but still experiences notable degradation in noisy conditions. Mixtral 8×22B emerges as the most vulnerable, exhibiting the lowest absolute performance and struggling particularly with exact answer generation.
@@ 5.3. Performance based on Different Error Types
In this section, we conduct an in-depth analysis of the impact of different types of OCR errors. Insertion, deletion, and substitution on QA systems. We use Gemma-2 (27B) for this analysis, as it was found to consistently outperform the other models across English, French, and German. As detailed in Section 4.1, each error type is categorized into three levels: Low, Medium, and High, where Low represents minimal presence and High indicates the most frequent occurrence of a particular type of error. We evaluated MultiOCR-QA’s performance across these categories, as illustrated in Figure 4.
@@
Deletion errors impact sentence coherence and factual consistency, especially when they corrupt or remove key words or essential contextual phrases. Although the impact is less pronounced at lower levels, it escalates sharply as the frequency of deletions increases. At higher levels of deletion error, the degradation in BERTScore is similar to that seen with substitution errors, highlighting how missing characters or words disrupt structured text.
@@
Summary of findings: Across languages, the results reveal that English and French exhibit similar degradation patterns, with BERTScore progressively decreasing as the OCR error frequency increases. However, in German, the sharpest decline is observed across all error types, particularly for substitutions and deletions. This suggests that OCR noise in German is more detrimental, probably because the older scripts have content in old German where characters such as (long s) are used instead of ”s”, which can be often misread as ”f” or ”l”.
@@
5.4. Performance based on Pre-Processed and Post-Processed OCR Text
In this section, we conduct an additional study to assess the impact of OCR on QA through pre-processing and post-processing Raw OCR text. In this experiment, we focus exclusively on paragraphs and their corresponding questions derived solely from the ICDAR dataset for the English language, comprising 83 paragraphs and 513 questions. We chose this subset of paragraphs and questions because its OCR ground truth is manually curated, offering a robust basis for comparison, unlike the CorrectedOCR text in the ChroniclingAmerica QA, which has been refined using GPT 3.5 (Brown et al., 2020). Our approach involves two strategies: pre-processing, where RawOCR text is corrected using the Gemma-2 27B model prior to answering the questions, and post-processing, where answers generated from RawOCR text are subsequently corrected. Gemma-2 27B was used in this experiment as it is identified as the best performing model among all other LLMs as demonstrated in Section 5.2. The results for each strategy are presented in Table 7.
Table 7. Performance Metrics of QA Systems Using Pre-Processed and Post-Processed OCR Text. Red indicates the percentage drop in performance when using different variants of paragraph used as context, compared to Corrected Paragraph (Ground Truth).
@@ 5.4. Performance based on Pre-Processed and Post-Processed OCR Text
Corrected Paragraph (Ground Truth) 63.99 58.09 16.76
@@
The results in Table 7 indicate that using ground truth paragraph as context gives better performance metrics, with a BERTScore of 63.99, Contains of 58.09, and EM of 16.76, compared to using RawOCR text or LLM corrected passage as context. The comparative analysis reveals that pre-processing the RawOCR text significantly enhances QA performance, as evidenced by higher BERTScore, Contains, and EM scores. This suggests that correcting OCR errors prior to QA processing preserves the semantic integrity of the text, thereby facilitating more accurate answer generation. However, it shows a percentage drop of 7. 28% compared to the Ground Truth BERTScore. In contrast, post-processing the generated answer is less effective, resulting in the highest drop of 12.33% and the lowest BertScore among others. This is likely because that approach does not include any contextual reference and leads to generation of corrected words that have limited relation to context.
@@
These findings emphasize the importance of integrating OCR error correction early in the QA pipeline to improve the reliability of QA systems, especially when dealing with historical texts or other archival materials. However, given the huge collections of digitized content with vastly varying levels of OCR quality that the current memory institutions (archives, libraries, museums, etc.) held, the correction cost and effort would be enormous. It is also difficult to correct the queried texts at inference time as this would also introduce computational cost and latency in online systems. Therefore, more robust QA systems that are aware of OCR errors and capable of predicting correct answers based on contextual information are required.
@@ 6. Conclusion
In this paper, we introduce MultiOCR-QA, a multilingual question-answering dataset for historical texts processed by OCR. Our dataset is unique as it is specifically designed to assess the impact of OCR errors on QA system performance, enabling a systematic analysis of how different error types—insertions, deletions, and substitutions- affect the accuracy and robustness of large language models. By incorporating both CorrectedOCR (clean text) and RawOCR (OCR-ed text), MultiOCR-QA facilitates comparative assessments that highlight the challenges posed by OCR noise in historical document processing. Our evaluation of multiple state-of-the-art LLMs demonstrates that OCR errors significantly degrade QA performance, with the most pronounced effects observed in languages with higher OCR error rates. Although larger models such as Gemma-2 27B and Qwen-2.5 72B exhibit greater resilience to OCR noise, smaller models and low-resource QA systems show a more substantial performance drop. These findings underscore the need for OCR-aware training strategies, particularly for historical document processing and multilingual QA tasks.
Use cases: The MultiOCR-QA dataset offers a unique resource to advance research on OCR-aware QA and studying QA on noisy OCR text, making it useful in several ways; It can be used to train LLMs to improve error correction capabilities and enhancing robustness against OCR inaccuracies while preserving the archaic language structure. It can be also used to expand LLMs’ multilingual processing abilities by training on multilingual OCR texts, improving accuracy in low-resourced languages.

### v2

@@ 4.1. Quantifying and Filtering Noise
Outlier Detection: To ensure a reliable analysis, we applied the Interquartile Range (IQR) method to detect and remove outliers in CER values. Outliers were defined as CER values below Q1 - 1.5 * IQR or above Q3 + 1.5 * IQR, where Q1 and Q3 represent the 33th and 66th percentiles of the CER distribution, respectively, and IQR is the difference between Q3 and Q1. This filtering resulted in the removal of 25 English, 351 French, and 2,423 German paragraphs.
@@
  • English: Low: CER << 0.0576, Medium: $0.0576\leq$ CER << 0.1238, High: CER $\geq$ 0.1238
  • French: Low: CER << 0.0357, Medium: $0.0357\leq$ CER << 0.0558, High: CER $\geq$ 0.0558
  • German: Low: CER << 0.2453, Medium: $0.2453\leq$ CER << 0.2757, High: CER $\geq$ 0.2757
@@
Mean Substitutions 54.97 31.51 240.23
Median Substitutions 38.00 21.00 230.50
Median Deletions 64.42 41.09 85.09
Median Deletions 18.00 25.00 39.00
Median Insertions 66.91 20.90 82.60
Median Insertions 27.50 11.00 80.00
@@ 5. Experiments and Results
In this section, we conduct a comprehensive analysis of MultiOCR-QA from several perspectives. First, we evaluate the performance of MultiOCR-QA across various LLMs, comparing different model families and sizes to assess their effectiveness in handling OCR-generated text. Second, we investigate the impact of OCR errors on QA performance, focusing on three main error types: insertion errors, deletion errors, and substitution errors.
@@ 5.2. Experimental Results
Tables 3, 4, and 5 present the impact of OCR errors on the performance of various LLMs in question-answering tasks for English, French, and German texts.
English text: Table 3 presents the performance of LLMs on English text using both CorrectedOCR (CP) and RawOCR (RP) paragraphs. Across all models, the transition from CP to RP negatively impacts performance. The best-performing model on CP is LLaMA-3.3-70B with a BERTScore of 66.94, followed by Gemma-2 27B at 63.99. When switching to RP, LLaMA-3.3-70B still achieves the highest BERTScore 62.87, but with a 6.08% drop, highlighting its robustness. The lowest-performing model is Mixtral 8×22B, showing the most significant impact of OCR errors. LLaMA-3.3-70B model achieves the best performance in CP for the Contains metric of 63.90, indicating its strong retrieval ability in clean text. However, it experiences a 21.3% drop in RP, suggesting moderate sensitivity to OCR errors. The lowest performing model in RP is again Mixtral 8×22B for the Contains metric, which sees a 15.70% decrease, indicating its greater vulnerability to noisy text. The Exact Match (EM) metric shows the steepest decline across models, emphasizing that OCR errors severely impact the models’ ability to generate precise answers.
French text: In Table 4 we summarize how LLMs perform on QA over French text. Models consistently perform better on CorrectedOCR, confirming and quantifying the negative influence of OCR errors on the QA accuracy. Gemma-2 27B model achieves the highest BERTScore of 76.51 on CP, demonstrating its strong ability to capture semantic similarity. Despite the 1.46% drop, it still maintains the highest performance with a BERTScore of 75.39 on RP, indicating robustness to OCR noise. Mixtral 8×22B model shows the smallest drop of 1.05%, but its overall score remains lower than the ones for the other models. For Contains, Qwen-2.5 72B achieves the highest score on CP 57.55, highlighting superior retrieval performance on clean text. However, it experiences a 19.90% drop in RP, reinforcing its vulnerability to OCR errors. LLLaMA-3.1 8B also struggles, with a 19.63% decline, showing difficulty in retrieving spans from noisy text. In terms of EM, Gemma-2 27B outperforms all models with an EM of 17.46 on CP. Although it drops by 15.23%, it still maintains the best EM value 14.80 in RP. Mixtral 8×22B model is the most affected in EM, dropping by 27.99%, suggesting that OCR noise drastically reduces its accuracy in generating exact answers.
@@
German text: Table 5 focusing on German language shows the most significant performance decline among the three languages. Unlike English and French, German exhibits the largest performance drop due to its lower OCR quality, and the models struggle more with its linguistic structure. Among the models evaluated, Gemma-2 27B achieves the highest BERTScore of 67.07 on CP, confirming its strong ability to capture semantic similarity in clean text. It also maintains the best performance on RP (63.78); however, it still experiences a 4.91% decrease, highlighting the adverse effects of OCR errors. In contrast, LLaMA-3.1 8B and LLaMA-3.3 70B show the biggest BERTScore drop (8.74% and 8.43% respectively), suggesting that these models struggle more with OCR noise. For the Contains metric, LLaMA-3.3 70B achieves the highest Contains score on CP, making it the most effective for retrieving relevant information in clean text. However, all models suffer a severe drop in Contains when moving to RP, with reductions exceeding 65%. Qwen-2.5 7B has the worst drop 66.40%, indicating that it faces challenges to retrieve information from noisy text. In terms of EM, Gemma-2 27B achieves the highest EM score on CP 2.69, while Mixtral 8×22B has the lowest EM 0.078. EM scores drop drastically across all models, the largest decrease is 79.04% for Gemma-2, reinforcing that word-level distortions from OCR errors make exact answer matching nearly impossible. The Mixtral model drops 72.74% in EM, making it highly unreliable for exact answers in noisy OCR text.
Summary of findings: OCR errors consistently degrade the performance of the models in English, French and German texts, resulting in maximum drop of 9.23%, 1.75%, and 8.74% in BERTScore, respectively. The most severe impact was observed in German due to the lower quality of the OCR and the complex linguistic structure. While larger models like Gemma-2 27B and LLaMA-3.3 70B demonstrate greater resilience, all models suffer substantial declines in Contains and Exact Match (EM) metrics, highlighting their weakness in retrieving and generating precise answers from noisy text. Gemma-2 27B consistently outperforms others, maintaining the highest BERTScore and EM across all languages, but still experiences notable degradation in noisy conditions. Mixtral 8×22B emerges as the most vulnerable, exhibiting the lowest absolute performance and struggling particularly with exact answer generation.
@@ 5.3. Performance based on Different Error Types
In this section, we conduct an in-depth analysis of the impact of different types of OCR errors: insertion, deletion, and substitution on QA systems. We use Gemma-2 (27B) for this analysis, as it was found to consistently outperform the other models across English, French, and German. As detailed in Section 4.1, each error type has been categorized into three levels: Low, Medium, and High, where Low represents minimal presence and High indicates the most frequent occurrence of a particular type of error. We evaluated MultiOCR-QA’s performance across these categories, as illustrated in Figure 4.
@@
Deletion errors impact sentence coherence and factual consistency, especially when they corrupt or remove key words or essential contextual phrases. Although the impact is less pronounced at lower levels, it escalates sharply when the frequency of deletions increases. At higher levels of deletion error, the degradation in BERTScore is similar to that seen with substitution errors, highlighting how missing characters or words disrupt structured text.
@@
Summary of findings: Across languages, the results reveal that English and French exhibit similar degradation patterns, with BERTScore progressively decreasing as the OCR error frequency increases. However, in German, the sharpest decline is observed across all error types, particularly for substitutions and deletions. This suggests that OCR noise in German is more detrimental, probably because the older scripts have content in old German where characters such as (long s) are used instead of ”s”, which can often be misread as ”f” or ”l”.
@@
Table 6. Performance metrics of QA systems on English text using pre-processed and post-processed OCR input. Red values indicate the percentage drop in performance compared to using Corrected Paragraph (Ground Truth) as context.
@@
Corrected Paragraph (Ground Truth) 63.65 57.60 16.20
@@
Table 7. Performance metrics of QA systems on French text using pre-processed and post-processed OCR input. Red values indicate the percentage drop in performance compared to using Corrected Paragraph (Ground Truth) as context.
Approach BERTScore Contains EM
Corrected Paragraph (Ground Truth) 65.17 52.40 14.40
[…]
Table 8. Performance metrics of QA systems on German text using pre-processed and post-processed OCR input. Red values indicate the percentage drop in performance compared to using Corrected Paragraph (Ground Truth) as context.
Approach BERTScore Contains EM
Corrected Paragraph (Ground Truth) 59.80 12.80 3.40
[…]
5.4. Effectiveness of Error Correction Strategies
Finally, we present an additional study to mitigate the impact of OCR errors on QA through two mitigation approaches, context correction and answer correction of Raw OCR text. We want to test simple solutions to reduce the effect of OCR noise in QA while preserving the syntactic and semantic integrity of the input.
Such solutions should also be feasible to be deployed in real-world settings with limited computational and financial resources.
We randomly sample 500 questions for each of the three languages. As the context correction approach [^11] The prompt used for context correction. System Prompt: You are an expert at understanding the historical texts and correcting OCR errors. User Prompt: You are provided with a historical English text containing spelling mistakes. Correct only the spelling mistakes and present the corrected text in a single paragraph. If you cannot correct the mistakes, reply with ”Not able to correct.” Historical Text : {context}, we correct the RawOCR text first using the Gemma-2 27B model and then use it as context for answering questions. In contrast, in the answer correction approach, we use RawOCR text as context to generate an answer first, and then we correct it afterward.
The context correction approach may improve performance by correcting the entire context before QA. However, it is computationally expensive due to the need to process the entire passage. On the other hand, the Answer correction approach is more lightweight and cost-efficient, as it only involves correcting short answer spans. Nevertheless, it may suffer from context misalignment, especially when the input is both noisy and lacks sufficient contextual information. We use Gemma-2 27B for both strategies as it demonstrated the strongest performance across all languages in our main experiments (Section 5.2). The results of this analysis are summarized for different languages in Table 6, Table 7, and Table 8, respectively. In the tables, the context correction approach is labeled as LLM Corrected Paragraph, while the answer correction approach is labeled as RawOCR Corrected Answer.
The results across English, French, and German (Tables 6, 7, and 8) consistently show that using corrected (ground-truth) paragraphs as context yields the highest QA performance across all metrics. Among the three languages, English shows the greatest benefit from pre-processing RawOCR with LLM correction, with BERTScore and EM metrics improving significantly over using RawOCR directly. In contrast, French and German show minimal or even negative gains from pre-processing, with German showing extremely low EM score with degradation of 94.12% compared to the ground truth. For all three languages, post-processing the answer (i.e., correcting the generated answer after QA) consistently results in the worst performance, suggesting that once the model has reasoned over noisy text, semantic distortions are difficult to recover. Across all languages, the post-processing strategy, where the answer is corrected after being generated from RawOCR input, consistently results in the lowest performance, highlighting that once a model has reasoned over noisy context, semantic errors become difficult to correct.
@@ 5.4. Effectiveness of Error Correction Strategies
These findings emphasize the importance of integrating OCR error correction early in the QA pipeline to improve the reliability of QA systems, especially when dealing with historical texts or other archival materials of lower quality. However, given the huge collections of digitized content with vastly varying levels of OCR quality that the current memory institutions (archives, libraries, museums, etc.) held, the correction cost and effort would be enormous. It is also difficult to correct the queried texts at inference time as this would also introduce computational cost and latency in online systems. Therefore, more robust QA systems that are aware of OCR errors and capable of predicting correct answers based on contextual information are required.
@@ 6. Conclusion
In this paper, we conducted a comprehensive evaluation of QA systems under the influence of OCR-induced noise using multilingual historical texts. To facilitate this analysis, we introduced MultiOCR-QA, the first multilingual QA dataset derived from OCR-processed historical documents in English, French, and German. The dataset includes both RawOCR (noisy) and CorrectedOCR (clean) text, enabling controlled and comparative assessments of model robustness under real-world digitization errors. By leveraging both CorrectedOCR and RawOCR we systematically analyzed how different types of OCR errors—insertions, deletions, and substitutions affect the performance of LLMS in QA.
Our experiments reveal that OCR noise leads to substantial degradation in QA performance across all languages and metrics, with particularly severe impacts in settings with high error rates. While larger LLMs such as Gemma-2 27B and Qwen-2.5 72B show relatively better robustness, even these models experience significant accuracy drops when processing noisy text. Moreover, pre- and post-processing strategies using LLMs to correct OCR errors offer only limited benefits and can, in some cases, degrade performance even further. These findings highlight the limitations of current QA systems in handling real-world noisy text and emphasize the need for evaluation frameworks and model designs that can handle OCR distortions, especially in historical and multilingual contexts.
Use cases of MultiOCR-QA: The MultiOCR-QA dataset offers a unique resource to advance research on OCR-aware QA and studying QA on noisy OCR text, making it useful in several ways. It can be used to train LLMs to improve error correction capabilities and enhancing robustness against OCR inaccuracies while preserving the archaic language structure. It can also be used to expand LLMs’ multilingual processing abilities by training on OCR text in multiple languages, thereby enhancing performance in languages beyond English.
@@
7. Usage of Generative AI
This work involved the use of Generative AI (GenAI) models at several stages of the research process. The MultiOCR-QA dataset has been automatically generated using instruction fine-tuned large language model (LLaMA 3.1-70b) for English, French, and German. The models were explicitly used for answer span extraction and question generation based on historical CorrectedOCR text. Various generative models, including LLaMA, Qwen, Mixtral, and Gemma, were used in the downstream evaluation of QA performance across corrected and noisy OCR text. Additionally, GenAI tools such as ChatGPT and Gemini were occasionally used to assist with resolving coding bugs or syntax errors in scripts or LaTeX syntax.
All conceptual work, dataset engineering, code design, and writing remain the sole intellectual contribution of the authors. GenAI tools were used in a supportive capacity, similar to that of debugging assistants or grammar checkers, without contributing novel research content.